OpenStudy (anonymous):

Medal + Fan Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example, can be written as 7sqrt(a).

OpenStudy (anonymous):

@bishopbabyy0362

OpenStudy (anonymous):

Thanks. Is that all?

OpenStudy (anonymous):

lol, you don't have to be sorry. I just wanted to make sure I didn't turn in my work half done. Thanks so much again!

OpenStudy (anonymous):

\[\sqrt{49x ^{3} } \ \ = \ \ \sqrt{7^{2} x ^{2} *x}= \ \ 7x \sqrt{x} \]